# Rehearsal-free Federated Domain-incremental Learning

### v1

@@ 4.1 Experimental Setup
However, they were originally proposed for centralized continual learning contexts, and consider to FDIL is an under-explored field. We adapted these methodologies for our FDIL setting, resulting in variants we refer to as FedLwF, FedEWC, FedL2P and FedDualPrompt, respectively. Particularly, to ensure a fair comparison, we initially deactivated the prompt pool feature in FedL2P and FedDualPrompt. Moreover, to demonstrate the superior efficacy of our RefFiL approach, even when competing methods employ prompt pool as a rehearsal technique, we also evaluate versions of FedL2p and FedDualPrompt with the prompt pool feature reactivated. These special versions are indicated with a superscript in the Table 1 3 2 4.
Evaluation Metrics. Our evaluation primarily includes top-1 accuracy, reflecting the model’s most probable correct prediction. We also report Average (Avg %) accuracy, as defined in iCaRL [30], which is the mean accuracy across all learning steps, indicating the model’s stability and consistency. Additionally, we present the Final (Last %) accuracy, measured after the last learning step, to assess the model’s knowledge retention and consolidation capabilities.
@@
Table 1: Summarised results on four datasets: Digits-five with five incremental tasks, FedDomainNet with six, and both OfficeCaltech10 and PACS with four tasks. Displaying average accuracy (Avg%) and final task accuracy (Last%). The amount of change in our method compared to other methods is marked as $\Delta$, The results of our RefFiL are highlighted in colour
@@
Table 2: Summarised results in new domain order of Table 1.
@@
Table 3: Comparison of RefFiL’s performance with five baseline methods on four widely used datasets, showcasing average accuracy (Avg %) and accuracy for each domain task (%). At each task step, all clients have an equal number of classes but quantity shift data.
@@
Table 4: Comparison of RefFiL’s performance in new domain order of table 3.
Implementation details. In our study, all methods are implemented using PyTorch [26] and run on a single NVIDIA RTX 4090 GPU, using ResNet10 [14] as the feature extractor classification model backbone. We designed a simple embedding model as the feature map tokenizer, similar to ViT [10], with initialized-only and frozen parameters for feature embedding. The FedAvg [24] algorithm was employed for global model aggregation. Each task involves training the model over $R=30$ rounds. The number of epochs for local updates on each client is set at 20. We use learning 0.04 for FedDomainNet, 0.06 for OfficeCaltech10 and 0.03 for the rest of the datasets to get the proper performance. Additionally, we use SGD as an optimizer for all experiments. Unless specified otherwise, the constraint factor $\lambda$ in the EWC method is fixed at 300, and the temperature for distillation is set to the default value of 2. Our client increment setting aligns with what we outlined in Appendix A. Specifically, for experiments on the Digit-Five, PACS, and FedDomainNet datasets, we start with 20 clients, select 10, and increment by 2 clients with each new task. Conversely, for experiments on the OfficeCaltech10 dataset, considering the data sample limitation, we start with 10 clients, select 5, and increment by 1 client per new task. In all experiments, 80% of the $\mathcal{U}o$ old clients transition to a new task for each respective task. For the hyperparameters in Eq. 7, we fix $\tau_{\text{min}}$ at 0.3, $\gamma$ at 0.1, $\beta$ at 0.05, and $\tau$ at 0.9.
4.2 Qualitative Results
@@
In our study, we evaluated the performance of seven baseline methods alongside our proposed method, RefFiL, across four widely used datasets: Digit-Five, OfficeCaltech10, FedDomainNet, and PACS. The results, detailed in Table 1, Table 3, Table 2, and Table 4, indicate that RefFiL consistently outperforms the baseline methods in both averaged and last task accuracy, even under the impact of catastrophic forgetting. Notably, RefFiL maintains its advanced performance consistently, regardless of changes in domain order. To provide a comprehensive comparison, we conducted three additional experiments, which are discussed in Appendix D. These experiments illustrate the accuracy distribution across different domains as tasks change and analyze the decision-making boundaries of all baseline methods.
@@
RefFiL efficiently mitigates forgetting. We report detailed performance in Table 3. The means the baseline enabled prompt pool to save generated prompts and then be sampled one for better mitigate forgetting. And the result can be summarised in Table 1, which demonstrates that RefFiL consistently achieves state-of-the-art (SOTA) results, showcasing robust performance across all four datasets in terms of average accuracy (Avg) and performance on the final task of each dataset (Last). Specifically, RefFiL shows a remarkable outperforming over the baseline methods in different scenarios. In smaller and less complex datasets like Digit-Five, RefFiL surpasses other baselines by 1.79% to 9.55% in average accuracy and by 1.94% to 16.22% in the final task accuracy. In the OfficeCaltech10 dataset, RefFiL outperforms baselines by 5.70% to 9.18% in average accuracy and 5.90% to 18.11% in the last task accuracy. Moreover, in larger and more complex image datasets like PACS, RefFiL continues to lead, achieving 1.27% to 15.20% higher average accuracy and 3.20% to 17.66% higher accuracy in the final task compared to the baselines. Similarly, despite the challenges posed by the FedDomainNet dataset, such as sparse data distribution across 48 classes and high-resolution images. RefFiL still demonstrated improvement. RefFiL outshines others with an improvement of 0.40% to 6.75% in average accuracy and 0.56% to 3.39% in the final task accuracy. These results collectively underscore the effectiveness and adaptability of RefFiL in handling diverse and challenging datasets, maintaining consistently high performance both overall and in the crucial final tasks.
RefFiL produces domain order-independent consistency results.
@@
To demonstrate the consistency of our results across different domain introduction orders, we generated new results presented in Table 4 and summarized in Table 2. In these experiments, we randomly altered the order of domain introductions. Specifically, in Table 4, the domain orders for Digit-Five, OfficeCaltech10, and FedDomainNet were significantly disrupted, while in PACS, only the order of the first two domains was changed.
This setup highlights the robustness of our method, showing that RefFiL maintains high performance despite variations in the sequence of domain introductions. Compared to the results in Table 1, although FedDualPrompt and FedDualPrompt achieved better performance in the Last metric, RefFiL consistently outperformed these methods in the more critical Avg metric, indicating its superior overall performance. For both the OfficeCaltech10 and PACS datasets, RefFiL’s advantages in Avg and Last metrics were magnified compared to most of the baselines. In the most challenging dataset, FedDomainNet, while RefFiL’s advantage in Avg reduced from 0.40% to 0.20%, its Last metric slightly improved from 0.93% to 1.20%. These results significantly amplified RefFiL’s advantages over other baselines.
4.3 Ablation Study
@@
Table 5: Ablation study of RefFil on OfficeCaltech10 dataset, and the amount of improvement in RefFiL with different components compared to baseline (Finetune) is marked as $\Delta$.
To understand how different components can mitigate catastrophic forgetting problems, we conduct the experiments as shown in Table 5; we dissect the impact of each module within RefFil, as well as their synergistic effects. Each experimental configuration involving a separate component is indicated by a checkmark (✓).
@@
5 Conclusion

### v2

@@ V-A Experimental Setup
However, they were originally proposed for centralized continual learning contexts, and consider to FDIL is an under-explored field. We adapted these methodologies for our FDIL setting, resulting in variants we refer to as FedLwF, FedEWC, FedL2P and FedDualPrompt, respectively. Particularly, to ensure a fair comparison, we initially deactivated the prompt pool feature in FedL2P and FedDualPrompt. Moreover, to demonstrate the superior efficacy of our RefFiL approach, even when competing methods employ prompt pool as a rehearsal technique, we also evaluate versions of FedL2p and FedDualPrompt with the prompt pool feature reactivated. These special versions are indicated with a superscript in the Table I III II IV. Evaluation Metrics. Our evaluation primarily includes top-1 accuracy, reflecting the model’s most probable correct prediction. We also report Average (Avg %) accuracy, as defined in iCaRL [23], which is the mean accuracy across all learning steps, indicating the model’s stability and consistency. Additionally, we present the Final (Last %) accuracy, measured after the last learning step, to assess the model’s knowledge retention and consolidation capabilities. Additionally, we consider Forgetting (FGT), which quantifies the average amount of information lost across tasks, and Backward Transfer (BwT), which evaluates how learning new tasks influences the performance of previously learned tasks.
@@
TABLE I: Summarised results on four datasets: Digits-five with five incremental tasks, FedDomainNet with six, and both OfficeCaltech10 and PACS with four tasks. Displaying average accuracy (Avg%) and final task accuracy (Last%). The amount of change in our method compared to other methods is marked as $\Delta$, The results of our RefFiL are highlighted in colour
@@
TABLE II: Summarised results in new domain order of Table I.
Implementation details. In our study, all methods are implemented using PyTorch and run on a single NVIDIA RTX 4090 GPU, using ResNet10 as the feature extractor classification model backbone. We designed a simple embedding model as the feature map tokenizer, similar to ViT [6], with initialized-only and frozen parameters for feature embedding. The FedAvg [18] algorithm was employed for global model aggregation. Each task involves training the model over $R=30$ rounds. The number of epochs for local updates on each client is set at 20. We use a learning rate of 0.04 for FedDomainNet, 0.06 for OfficeCaltech10 and 0.03 for the rest of the datasets to get the proper performance. Additionally, we use SGD as an optimizer for all experiments. Unless otherwise specified, the constraint factor $\lambda$ in the EWC method is fixed at 300, and the temperature for distillation is set to the default value of 2. For the experiments on the Digit-Five, PACS, and FedDomainNet datasets, as shown in Tables I, II, III, IV, and VIII, the setup starts with 20 clients, of which 10 are initially selected, with 2 additional clients incrementally introduced per new task. For the OfficeCaltech10 dataset, considering its data sample limitations, the setup begins with 10 clients, of which 5 are initially selected, and 1 additional client is incrementally introduced per new task. Since this work focuses on domain-incremental learning, after the first task, all clients retain the same number of classes but represent distinct domains, and for each task $t$, 80% of the $M$ clients from task $t$ transition to the next task. In the experiments summarized in Table VI, the setup includes 10 clients initialized at the start, with 1 additional client joining per task increment, while for the results in Table V, the Digits Dataset is starting with 20 clients and incrementing by 2 clients per task. The hyperparameters in Eq. 10 are fixed as follows: $\tau_{\text{min}}=0.3$, $\gamma=0.1$, $\beta=0.05$, and $\tau=0.9$.
@@
TABLE III: Comparison of RefFiL’s performance with five baseline methods on four widely used datasets, showcasing average accuracy (Avg %) and accuracy for each domain task (%). At each task step, all clients have an equal number of classes but quantity shift data.
@@
TABLE IV: Comparison of RefFiL’s performance in new domain order of table III.
Sel 8, 80% of $M$ Sel 2, 80% of $M$ Sel 5, 50% of $M$ Sel 5, 90% of $M$
Method AVG Last FGT BwT AVG Last FGT BwT AVG Last FGT BwT AVG Last FGT BwT
Finetune 47.01 20.28 0.348 -0.290 47.06 25.79 0.287 -0.288 48.24 25.59 0.294 -0.252 47.39 25.00 0.313 -0.267
FedLwF 46.19 23.23 0.342 -0.286 49.22 29.53 0.262 -0.202 48.04 26.18 0.281 -0.228 47.81 26.38 0.288 -0.227
FedEWC 45.15 18.50 0.326 -0.259 45.26 25.20 0.234 -0.197 47.83 25.00 0.292 -0.253 47.64 22.24 0.315 -0.255
FedL2P 50.04 23.03 0.301 -0.222 48.50 29.53 0.265 -0.215 51.20 28.74 0.291 -0.232 51.41 27.36 0.311 -0.231
FedL2P 50.93 25.20 0.323 -0.243 47.76 28.15 0.259 -0.209 53.38 29.13 0.282 -0.211 51.86 27.95 0.301 -0.234
FedDualPrompt 50.45 27.95 0.296 -0.232 47.04 25.98 0.263 -0.199 50.65 28.74 0.307 -0.237 50.44 29.33 0.305 -0.246
FedDualPrompt 50.06 25.00 0.321 -0.252 47.53 27.95 0.279 -0.217 51.97 27.56 0.309 -0.231 52.20 29.13 0.295 -0.219
RefFiL 54.59 33.86 0.278 -0.198 49.22 30.12 0.237 -0.192 54.74 34.25 0.270 -0.198 54.22 31.89 0.274 -0.200
TABLE V: Results on the OfficeCaltech10 Dataset, following the same domain order as in Table III.
Sel 10, 90% of $M$
Method AVG Last FGT BwT
Finetune 79.80 53.49 0.405 -0.256
FedLwF 78.64 53.51 0.443 -0.276
FedEWC 78.30 51.98 0.317 -0.226
FedL2P 84.00 57.03 0.374 -0.214
FedL2P 85.42 58.88 0.323 -0.191
FedDualPrompt 84.57 58.06 0.331 -0.201
FedDualPrompt 84.70 57.71 0.324 -0.193
RefFiL (Our) 86.84 62.50 0.313 -0.182
TABLE VI: Results on the Digits Dataset, following the same domain order as in Table III.
V-B Qualitative Results
@@
In our study, we evaluated the performance of seven baseline methods alongside our proposed method, RefFiL, across four widely used datasets: Digit-Five, OfficeCaltech10, FedDomainNet, and PACS. The results, detailed in Table I, Table III, Table II, and Table IV, indicate that RefFiL consistently outperforms the baseline methods in both average and last task accuracy, even under the impact of catastrophic forgetting. Notably, RefFiL maintains its advanced performance, regardless of changes in domain order. Furthermore, to provide a comprehensive comparison across different settings and evaluation metrics, we conducted three additional experiments, detailed in Tables VI and V, which further exhibit the outstanding performance of RefFiL.
@@
RefFiL efficiently mitigates forgetting. We report detailed performance in Table III. The indicates baselines enabled with a prompt pool to save generated prompts, which are sampled to better mitigate forgetting. The summarized results in Table I demonstrate that RefFiL consistently achieves state-of-the-art (SOTA) performance across all four datasets in terms of both average accuracy (Avg) and final task accuracy (Last). For example, in smaller and less complex datasets such as Digit-Five, RefFiL significantly outperforms other baselines, achieving improvements in Avg by up to 9.55% and in Last by up to 16.22%. Similarly, in OfficeCaltech10, RefFiL exceeds baselines by as much as 9.18% in Avg and 18.11% in Last, showcasing its ability to effectively manage diverse domains. In larger, more complex datasets like PACS, RefFiL maintains its leadership, achieving Avg and Last improvements of up to 15.20% and 17.66%, respectively. Even under the challenging conditions of FedDomainNet, characterized by sparse data distributions across 48 classes, RefFiL demonstrates robust performance, improving Avg by up to 6.75% and Last by up to 3.39% compared to baselines.
To further evaluate RefFiL’s adaptability under diverse configurations, we conducted additional experiments, as detailed in Table V for OfficeCaltech10 and Table VI for Digit-Five. These experiments explore RefFiL’s performance under varying client selection percentages and task transfer configurations. In OfficeCaltech10, RefFiL achieves consistently superior results, with Avg reaching 54.59% and Last 33.86% under the select 8 clients, 80% task transfer setup. RefFiL also demonstrates the lowest FGT at 0.278 and the most favorable BwT at -0.198 in this setup, reflecting its exceptional capacity to retain knowledge from earlier tasks while effectively incorporating new learning. Compared to baselines such as FedLwF (FGT: 0.342, BwT: -0.286) and FedL2P (FGT: 0.301, BwT: -0.222), RefFiL significantly reduces forgetting and enhances backward transfer, achieving a more optimal balance between forward and backward learning. These results validate RefFiL’s ability to mitigate catastrophic forgetting while ensuring stable performance across diverse configurations.
Similarly, in Digit-Five, RefFiL effectively adapts to domain selection sizes, achieving 86.84% Avg and 62.50% Last in the select 10 clients, 90% task transfer setup. It maintains competitive forgetting (FGT: 0.313) and achieves the most favorable backward transfer (BwT: -0.182), further reinforcing its capacity to integrate new knowledge while minimizing performance degradation on prior tasks. These results underscore RefFiL’s adaptability and robustness across different scenarios, highlighting its ability to consistently deliver high performance and maintain stability in diverse FDIL tasks.
@@
RefFiL produces domain order-independent consistency results. To demonstrate the consistency of our results across different domain introduction orders, we generated new results presented in Table IV and summarized in Table II. In these experiments, we randomly altered the order of domain introductions. Specifically, in Table IV, the domain orders for Digit-Five, OfficeCaltech10, and FedDomainNet were significantly disrupted, while in PACS, only the order of the first two domains was changed.
@@
TABLE VII: Ablation study of RefFiL on OfficeCaltech10 dataset, and the amount of improvement in RefFiL with different components compared to baseline (Finetune) is marked as $\Delta$.
This setup highlights the robustness of our method, showing that RefFiL maintains high performance despite variations in the sequence of domain introductions. Compared to the results in Table I, although FedDualPrompt and FedDualPrompt achieved better performance in the Last metric, RefFiL consistently outperformed these methods in the more critical Avg metric, indicating its superior overall performance. For both the OfficeCaltech10 and PACS datasets, RefFiL’s advantages in Avg and Last metrics were magnified compared to most of the baselines. In the most challenging dataset, FedDomainNet, while RefFiL’s advantage in Avg reduced from 0.40% to 0.20%, its Last metric slightly improved from 0.93% to 1.20%. These results significantly boosted RefFiL’s advantages over others.
V-C Ablation Study
To understand how different components can mitigate catastrophic forgetting problems, we conduct the experiments as shown in Table VII; we dissect the impact of each module within RefFiL, as well as their synergistic effects. Each experimental configuration involving a separate component is indicated by a checkmark (✓).
@@ V-C Ablation Study
An important observation is that while combining CDAP with GPL can improve the model’s average performance, it does not outperform the individual use of CDAP or GPL. This may be due to the increased domain data heterogeneity in the final task. The personalized prompts generated by CDAP and shared by GPL might cause difficulty for the model in distinguishing between heterogeneous domain information.
@@
VI Hyperparameter sensitivity analysis
Exp $\tau$ $\tau_{\text{min}}$ $\gamma$ $\beta$ $\tau^{\prime}$ (3rd) Avg Last
1 0.5 0.2 0.15 0.1 0.325 42.17 36.22
2 0.5 0.4 0.05 0.05 0.425 42.32 37.01
3 0.7 0.3 0.1 0.05 0.560 43.14 37.80
4 0.9 0.2 0.05 0.1 0.675 42.89 36.81
5 0.9 0.4 0.05 0.01 0.837 42.30 36.02
w/o $\tau^{\prime}$ 41.97 34.39
Our 0.9 0.3 0.1 0.05 0.720 44.33 38.39
TABLE VIII: Experimental configurations for sensitivity analysis of hyperparameters in OfficeCaltech10 dataset, with domain order shown in Table IV.
To evaluate the sensitivity of the hyperparameters in Eq.10, we conducted a grid search. Based on our preliminary experiments, we identified a suitable range for each hyperparameter as follows: $\tau$ in ${0.5,0.7,0.9}$, $\tau_{\text{min}}$ in ${0.2,0.3,0.4}$, $\gamma$ in ${0.05,0.1,0.15}$, and $\beta$ in ${0.01,0.05,0.1}$. These parameters can form a total of 81 combinations. Given that the final value of $\tau^{\prime}$ influences the loss function $\mathcal{L}_{DPCL}$, we selected five representative groups of hyperparameter combinations to explore how these variations affect the model’s performance. Additionally, we conducted a comparative experiment without applying the overall decay parameter $\tau^{\prime}$, marked as w/o $\tau^{\prime}$. As illustrated in TableVIII, different combinations of these parameters resulted in only minor variations in the model’s performance metrics. Notably, in the last task, where the overall domain diversity increased, $\tau^{\prime}$ effectively mitigated the impact of heterogeneous domain data on model performance.
VII Conclusion